# Unsupervised Video Domain Adaptation with Masked Pre-Training and Collaborative Self-Training

### 05 Dec 2023 v1

@@ 3.2 Self-Supervised Initialization
In contrast to some other video domain adaptation works [10, 65] that use network weights from Kinetics-400 supervised pre-training, we initialize our network from self-supervised pre-training. As contended in [43] for UDA with images, we argue that supervised pre-training can potentially complicate the study of VUDA techniques. When the pre-trained network has been trained to classify categories present in the DA dataset, some DA techniques could perform well simply by preserving the capabilities of the pre-trained model despite not generalizing well to DA tasks that do not have category overlap with the pre-training dataset. Moreover, the UDA problem formulation implies that labeled instances of relevant classes are only available from the source domain. This condition does not hold for the Daily-DA and Sports-DA benchmarks that we use for evaluation, as 6 out of their 8 classes also appear in the Kinetics-400 dataset. Further exacerbating the issue with supervised Kinetics pre-training is the fact that Kinetics serves as a target domain in both benchmarks. Thus, we initialize the network weights from UMT [30] single-modality (i.e., video only) self-supervised pre-training on Kinetics-710 [29]. Please refer to the Appendix for select results using a UMT pre-trained network with additional supervised fine-tuning on Kinetics-400, which unsurprisingly achieves better baseline performance.
@@ 4.1 UNITE
Following the original UMT method, we leverage a CLIP [44] ViT-B image encoder as . We keep the standard UMT masking ratio of and perform alignment between the last 6 layers of the networks. is initialized as described in Sec. 3.2, along with the decoders , which are kept frozen throughout Stage 1.
@@
Stage 3: Collaborative Self-Training (CST). After enhancing the ability of to extract meaningful features in the target domain (Stage 1) and fitting for the classification task using source videos (Stage 2), we further adapt to target data using a self-training process. In a conventional self-training framework (denoted as in Eq. 2), the overall loss is a sum of losses on labeled source domain samples and select target domain samples, chosen using mask . The target domain loss leverages the model’s own predictions (where and denotes the softmax function) in lieu of labels in the CE loss computation. Various strategies have been used to form the selection mask , with the goal of choosing target samples whose predictions are more likely to be correct [48, 68, 43].
@@
The complete loss function for the collaborative self-training stage (depicted in Figure 3) can be expressed as:
@@ 5.2 Implementation Details
During Stage 1, we perform UMT training for 50 epochs on the target dataset using a batch size of 256 and a learning rate of 1.5e-5. During Stage 2, we use a batch size of 28, a learning rate of 2.5e-5, and apply a layer-wise learning rate decay of 0.65. During stage 3, we use a batch size of 40 (20 videos from each domain), a learning rate of 1e-5, a MatchOrConf confidence threshold of and a target domain loss weight . Training in all stages uses the AdamW optimizer [35] and a weight decay of 0.05.
Please refer to the Appendix for more information about hyperparameters and augmentations used in each stage.
Zero-Shot Classification with CLIP. In Stage 3 of UNITE, we use CLIP (ViT-B/16) in a zero-shot manner to refine the pseudolabels of target domain videos. We do so by passing each of the video frames through the CLIP image encoder and computing the cosine similarity with a set of text representations produced by the CLIP text encoder. The inputs to the text encoder are the class names associated with the particular action recognition task, embedded in the following template: “A video of a person {class}.” To form a video-level prediction, we simply average the softmax of the cosine similarity scores across the video frames. Refer to the Appendix for a listing of class names used for each task.
@@ 5.4 Benchmark Results
Tables 1, 2 and 3 show the results of our evaluation of UNITE on Daily-DA, Sports-DA and UCFHMDB. We include accuracies after Stage 2 of the pipeline (denoted as UNITE w/o CST) as well as after applying the full UNITE process. For reference, we include accuracies using our UMT pre-trained network with source domain supervised fine-tuning (Source Only) and with target domain supervised fine-tuning (Target Only). We also incluide the lower and upper bounds (denoted in the tables as LB and UB) of DALL-V [67] (a CLIP-based ResNet-50 model) for added context. We additionally display the zero-shot classification accuracy of the CLIP teacher model that we utilize during Stage 1 and Stage 3 of UNITE.
@@ 6 Additional Analysis & Discussion
Student exceeds the teacher. We see in the benchmark results that the video student model exceeds the zero-shot accuracy of the CLIP teacher in virtually all domain shifts. This suggests that UNITE successfully leverages the spatial modeling capabilities of CLIP to train an even more powerful spatiotemporal model for the action recognition task. We can see from Figure 4 that the teacher model and the student model prior to collaborative self-training have differing capabilities. For example, the image teacher is stronger at recognizing actions from the jump class than the Stage 2 student. On the other hand, the video student is significantly better than the teacher at recognizing wave at the start of CST. In both cases, we see that the student model, through collaborative self-training, comes to exceeds both the Stage 2 student and the CLIP image teacher. With the MatchOrConf pseudolabeling scheme, the two models are able to work together to achieve higher accuracy in the target domain.
Impact of Stage 2 and 3. In Table 4 we ablate the stages in UNITE. While we observe that UMT pre-training and collaborative self-training each offer a benefit in the absence of the other, the largest gains in domain transfer performance are obtained when combining the two stages. For example, on the ARIDHMDB domain shift from Daily-DA, Stage 1 and Stage 3 alone improve upon source only accuracy by +1.8% and +11.2% respectively. Meanwhile, the full UNITE pipeline results in +18.3% accuracy improvement on the target domain, suggesting that the UMT pre-training and collaborative self-training processes complement each other.
@@
Impact of data domains during pre-training. In Table 5, we assess the impact of utilizing each of the two data domains during UMT pre-training on UDA performance. While we find the results to vary across domain shifts, we observe the most consistent improvements when pre-training only on target domain videos. Although incorporating both domains during unsupervised pre-training has been successfully applied in image-based domain adaptation [46, 43, 27, 37], we find it to be suboptimal with UMT training on the VUDA tasks studied in this work.
@@
Comparison of pseudolabeling strategies. In Table 7, we explore the use of alternative pseudolabeling strategies to the MatchOrConf [69] scheme that we employ in UNITE. The first 3 rows of the table assess variations of the masked consistency pseudolabeling scheme proposed in PACMAC [43], where target samples are selected if there is agreement between predictions of masked views. This criterion can also be combined with a constraint on the confidence exceeding threshold . While we find the combination of these two constraints (ConsOrConf) to be effective for self-training, we find that MatchOrConf results in more consistent target accuracy improvements across domain shifts. We also attempt to combine MatchOrConf with a masked consistency constraint as a way to reduce the rate of errors where the student and teacher models agree on an incorrect prediction. This scheme, denoted as MatchAndConsOrConf adds an additional constraint to the matching criterion by also requiring agreement between the student model predictions of masked views. We find that this approach underperforms compared to the vanilla MatchOrConf scheme due to low utilization of target domain samples.
@@ 8 Additional Training Details
In Tables 10, 11 and 12 we provide a more detailed account of the training configurations used in each of the three stages of UNITE.
@@ 11 Zero-Shot Classification with CLIP
The process we use to perform zero-shot image classification using CLIP [44] is described in Sec. 5.2. In Table 15, we provide the exact class names used to form the inputs to the CLIP text encoder. Classification is performed using a single template: “A video of a person {class}.”

### 21 Mar 2024 v3

@@ 3.2 Self-Supervised Initialization
In contrast to some other video domain adaptation works [10, 65] that use network weights from Kinetics-400 supervised pre-training, we initialize our network from self-supervised pre-training. As contended in [43] for UDA with images, we argue that supervised pre-training can potentially complicate the study of VUDA techniques. When the pre-trained network has been trained to classify categories present in the DA dataset, some DA techniques could perform well simply by preserving the capabilities of the pre-trained model despite not generalizing well to DA tasks that do not have category overlap with the pre-training dataset. Moreover, the UDA problem formulation implies that labeled instances of relevant classes are only available from the source domain. This condition does not hold for the Daily-DA and Sports-DA benchmarks that we use for evaluation, as 6 out of their 8 classes also appear in the Kinetics-400 dataset. Further exacerbating the issue with supervised Kinetics pre-training is the fact that Kinetics serves as a target domain in both benchmarks. Thus, we initialize the network weights from UMT [30] single-modality (i.e., video only) self-supervised pre-training on Kinetics-710 [29]. Please refer to Sec. 10 in the Appendix for select results using a UMT pre-trained network with additional supervised fine-tuning on Kinetics-400, which unsurprisingly achieves better baseline performance.
@@ 4.1 UNITE
Following the original UMT method, we leverage a CLIP [44] ViT-B image encoder as . We keep the standard UMT masking ratio of and perform alignment between the last 6 layers of the networks. is initialized as described in Sec. 3.2, along with the projections , which are kept frozen throughout Stage 1.
@@
Stage 3: Collaborative Self-Training (CST). After enhancing the ability of to extract meaningful features in the target domain (Stage 1) and fitting for the classification task using source videos (Stage 2), we further adapt to target data using a self-training process (depicted in Figure 3). In a conventional self-training framework (denoted as in Eq. 2), the overall loss is a sum of losses on labeled source domain samples and select target domain samples, chosen using mask . The target domain loss leverages the model’s own predictions (where and denotes the softmax function) in lieu of labels in the CE loss computation. Various strategies have been used to form the selection mask , with the goal of choosing target samples whose predictions are more likely to be correct [48, 68, 43].
@@
The complete loss function for the collaborative self-training stage can be expressed as:
@@ 5.2 Implementation Details
During Stage 1, we perform UMT training for 50 epochs on the target dataset using a batch size of 256 and a learning rate of 1.5e-5. During Stage 2, we use a batch size of 28, a learning rate of 2.5e-5, and apply a layer-wise learning rate decay of 0.65. During stage 3, we use a batch size of 40 (20 videos from each domain), a learning rate of 1e-5, a MatchOrConf confidence threshold of and a target domain loss weight . Training in all stages uses the AdamW optimizer [35] and a weight decay of 0.05. Please refer to Sec. 8 in the Appendix for more training details for each stage.
Zero-Shot Classification with CLIP. In Stage 3 of UNITE, we use CLIP (ViT-B/16) in a zero-shot manner to refine the pseudolabels of target domain videos. We do so by passing each of the video frames through the CLIP image encoder and computing the cosine similarity with a set of text representations produced by the CLIP text encoder. The inputs to the text encoder are the class names associated with the particular action recognition task, embedded in the following template: “A video of a person {class}." To form a video-level prediction, we simply average the softmax of the cosine similarity scores across the video frames. Refer to the Appendix (Sec. 11) for a listing of class names used for each task.
@@ 5.4 Benchmark Results
Tables 1, 2 and 3 show the results of our evaluation of UNITE on Daily-DA, Sports-DA and UCFHMDB. We include accuracies after Stage 2 of the pipeline (denoted as UNITE w/o CST) as well as after applying the full UNITE process. For reference, we include accuracies using our UMT pre-trained network with source domain supervised fine-tuning (Source Only) and with target domain supervised fine-tuning (Target Only). We also include the source only and target only accuracies (denoted in the tables as LB and UB) of DALL-V [67] (a CLIP-based ResNet-50 model) for added context. We additionally display the zero-shot classification accuracy of the CLIP teacher model that we utilize during Stage 1 and Stage 3 of UNITE.
@@ 6 Additional Analysis & Discussion
Student exceeds the teacher. We see in the benchmark results that the video student model exceeds the zero-shot accuracy of the CLIP teacher in virtually all domain shifts. This suggests that UNITE successfully leverages the spatial modeling capabilities of CLIP to train an even more powerful spatiotemporal model for the action recognition task. We can see from Figure 4 that the teacher model and the student model prior to collaborative self-training have differing capabilities. For example, the image teacher is stronger at recognizing actions from the jump class than the Stage 2 student. On the other hand, the video student is significantly better than the teacher at recognizing wave at the start of CST. In both cases, we see that the student model, through collaborative self-training, comes to exceeds both the Stage 2 student and the CLIP image teacher. With the MatchOrConf pseudolabeling scheme, the two models are able to work together to achieve higher accuracy in the target domain. Accuracies on the pick and run classes decline after CST, as self-training tends to reinforce the accuracy of well-performing classes, sometimes at the expense of more difficult ones that are infrequently assigned pseudolabels. Modifying CST to account for the learning status of each class, like in [68], may help remedy this issue.
Impact of Stage 1 and 3. In Table 4 we ablate the stages in UNITE. While UMT pre-training and collaborative self-training each offer a benefit in the absence of the other, the largest gains in domain transfer performance are obtained when combining the two stages. For example, on the ARIDHMDB domain shift in Daily-DA, Stage 1 and Stage 3 alone improve upon source only accuracy by +1.8% and +11.2% respectively. Meanwhile, the full UNITE pipeline results in +18.3% accuracy improvement on the target domain, suggesting that the masked pre-training and masked self-training processes complement one another.
@@
Impact of data domains during pre-training. In Table 5, we assess the impact of utilizing each of the two data domains during UMT pre-training on UDA performance. While the results vary across domain shifts, we observe the most consistent improvements when pre-training only on target domain videos. Although incorporating both domains during unsupervised pre-training has been successfully applied in image-based domain adaptation [46, 43, 27, 37], we find it to be suboptimal with UMT training on the VUDA tasks studied in this work.
@@
Comparison of pseudolabeling strategies. In Table 7, we explore the use of alternative pseudolabeling strategies to the MatchOrConf [69] scheme that we employ in UNITE. The first 3 rows of the table assess variations of the masked consistency pseudolabeling scheme proposed in PACMAC [43], where target samples are selected if there is agreement between predictions of masked views. This criterion can also be combined with a constraint on the confidence exceeding threshold . While we find the combination of these two constraints (ConsOrConf) to be effective for self-training, we find that MatchOrConf results in more consistent target accuracy improvements across domain shifts. We also consider combining MatchOrConf with a masked consistency constraint as a way to reduce the rate of errors where the student and teacher models agree on an incorrect prediction. This scheme, denoted as MatchAndConsOrConf adds an additional constraint to the matching criterion by also requiring agreement between the student model predictions of masked views. We find that this approach underperforms compared to the vanilla MatchOrConf scheme due to low utilization of target domain samples.
@@ 7 Conclusions
Acknowledgements. This research was sponsored by the Army Research Laboratory and was accomplished under Cooperative Agreement Number W911NF-21-2-0211. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ 8 Additional Training Details
In Tables 10, 11 and 12 we provide a more detailed account of the training configurations used in each of the three stages of UNITE. The ViT-B network we use has 87M parameters, all of which are trained in each of the 3 stages of UNITE. Stage 2, which is analogous to standard fine-tuning, took roughly 6 hours for KS (the largest task we evaluate on) whereas our additional stages, Stage 1 and Stage 3, took 1.5 hours and 6 hours respectively on 4 NVIDIA A5000 GPUs.
@@ 11 Zero-Shot Classification with CLIP
The process we use to perform zero-shot image classification using CLIP [44] is described in Sec. 5.2. In Table 15, we provide the exact class names used to form the inputs to the CLIP text encoder. Classification is performed using a single template: “A video of a person {class}."